OpenStudy (anonymous):

Setting up an algebraic equation, no "math" :-) Berta is planning a party for her friends. The room rental cost is $50 plus an additional $3 per person to cover food costs. Berta has saved up $100 dollars. What is the maximum number of people that can attend her party?

OpenStudy (igreen):

The room is $50. And an addition $3 per person. We can use the equation: 3x + 50 = 100, where x is the number of people.

OpenStudy (anonymous):

Oh okay awesome I know how to do I just didn't know where x went thanks so much :-)

OpenStudy (igreen):

To find out what the maximum number of people that can attend her party is, solve the equation I gave you. Can you do that?

OpenStudy (anonymous):

Yes I sure can. Is the equation the same thing in different form like this just wondering? 50+3x=100?

OpenStudy (igreen):

Yes, it's the same if you put it as 50 + 3x = 100. It'll be the same if you change their positions for Addition and Multiplication, but not for subtraction or division.

OpenStudy (anonymous):

Oh got it thank you again!

OpenStudy (igreen):

No problem.